OpenStudy (anonymous):

HELP!!! Find the curvature of r=2+2sin(t) at t=5pi/4

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Curvature#Curvature_of_a_graph use the formula for polar form

OpenStudy (dumbcow):

the derivatives aren't bad.... derivative of sin is cos , derivative of cos is -sin

OpenStudy (anonymous):

so I can use r as y equation?

OpenStudy (dumbcow):

no its different for polar equation....look below it gives the polar formula \kappa(\theta) = \frac{|r^2 + 2r'^2 - r r''|}{\left(r^2+r'^2 \right)^{3/2}}

OpenStudy (dumbcow):

\[\kappa(\theta) = \frac{|r^2 + 2r'^2 - r r''|}{\left(r^2+r'^2 \right)^{3/2}}\]

OpenStudy (anonymous):

Do I plug in 5pi/4 to get r, r', and r''?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

Ok but first I have to do the derivatives before I plug it in right?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Ok let me see if I can figure it, do you know the answer if I ask?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

Ok give me a sec

OpenStudy (anonymous):

Um.... is it 0.9799?

OpenStudy (dumbcow):

i get .135

OpenStudy (anonymous):

Ok so for r'=2cos(t) and r''=-2sin(t) right?

OpenStudy (dumbcow):

r = 2 - sqrt(2) r' = -sqrt(2) r'' = sqrt(2)

OpenStudy (dumbcow):

yes you are right

OpenStudy (anonymous):

Ok let me redo this

OpenStudy (dumbcow):

hold on, i apologize I made a mistake answer is .9799

OpenStudy (dumbcow):

i did by hand getting an exact answer first and i went too fast lol

OpenStudy (anonymous):

When I redid it I got .199, so my first answer was right?

OpenStudy (dumbcow):

hmm what did you change?

OpenStudy (anonymous):

I just plugged in sqrt2, 2-sqrt2 and -sqrt2. Not the actual decimal.

OpenStudy (dumbcow):

well the answers should be very close, .199 is way off from .97 check how you entered them into calculator yes i believe .9799 is correct

OpenStudy (anonymous):

Ok I'll redo it again I hate getting different answers lol

OpenStudy (anonymous):

Ok I got it thanks for the help :)

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

Find the parametric equations of the tangent lines to C at t = pi/6, curve C is r(t) = (4sin(3t))I+(2t)j+(4cos(3t))k.

OpenStudy (dumbcow):

ok so we need 3 lines, x(t) y(t) z(t) do each component the same way, the slope of tangent line is derivative at point pi/6 equation of line \[x(t) - r_x(\pi/6) = r_x'(\pi/6) (t - \pi/6)\] similar for y,z as well

OpenStudy (anonymous):

So would I set it up like this? \[4\sin(t)-r(\pi/6)=r'(\pi/6)(t-\pi/6)\]

OpenStudy (anonymous):

Oops I mean 4sin(3t)

OpenStudy (dumbcow):

almost, 4sin(3t) = r(t) for x component

OpenStudy (dumbcow):

x(t) is the dependent variable, leave it alone

OpenStudy (anonymous):

Oh so I would plug pi/6 into the 4sin(3t)?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

So to find y(t) and z(t) I would just replace the x(t) right?

OpenStudy (dumbcow):

yes the general form stays the same also change the r(t) and r'(t) to match the component

OpenStudy (anonymous):

Ok which is r(t)=4sin(3t) and r'(t)=12cos(3t)?

OpenStudy (dumbcow):

yes, plug in pi/6

OpenStudy (anonymous):

Is it x(t)=4?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Ah so now I just do y(t) and z(t) the same way?

OpenStudy (dumbcow):

yep, so find derivative of 2t to get r'(t) for y component

OpenStudy (anonymous):

If r'(t)=2 then how do I plug in pi/6?

OpenStudy (dumbcow):

you cant, no need, the slope is just 2

OpenStudy (anonymous):

Oh so y(t)=2t?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Ok let me do the z(t) and see if its right, one min

OpenStudy (anonymous):

Is it z(t)=-12t+2pi?

OpenStudy (dumbcow):

you got it

OpenStudy (anonymous):

YAY thank you so much!! I really appreciate it!

OpenStudy (dumbcow):

yw